A family friend has asked your help in analyzing the operations of three anonymous companies operating in the same service sector industry. Supply the missing data in the table below: (Loss amounts should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to nearest whole number.)

Company
A B C
Sales $430,000 $730,000 $600,000
Net operating income $54,000
Average operating assets $167,000 $145,000
Return on investment (ROI) 23 % 21 % %
Minimum required rate of return:
Percentage 15 % % 9 %
Dollar amount $54,000
Residual income $8,000
Answer #1
Company
A B C
Sales $430,000 $730,000 $600,000
Net operating income 167000*23% = 38410 $54,000 21050
Average operating assets $167,000 54000*100/21 = 257143 $145,000
Return on investment (ROI) 23 % 21 % 21050/145000 = 15 %
Minimum required rate of return:
Percentage 15 % 21% % 9 %
Dollar amount 167000*15% = 25050 $54,000 145000*9% = 13050
Residual income 38410-25050 = 13360 54000-54000 = 0 8000